Part C In Millikan's experiment, the charge on each drop of oil was measured in coulombs. Imagine the same experiment, but with charges measured in a fictitious unit called a zeet ( Z ). Drop Charge ( Z ) A 3.42×10−14 B 5.70×10−14 C 6.27×10−14 D 6.84×10−14 E 9.12×10−14 What is the charge on an electron in zeets?
Answer #1

: The reasoning involved in solving this question is that the charge on the oil drops must be derived from a integral number of excess electrons residing on the drops, since electrons bear a specific charge and are indivisible. We therefore look at the numbers representing the charges in zeets, and try to find the largest common factor. By inspection, the numbers 6.27 and 6.84 are the closest together, with a difference of 0.57, so this figure might be the required factor. Test this as follows:

3.42/0.57 = 6.0
5.70/0.57 = 10.0
6.27/0.57 = 11.0
6.84/0.57 = 12.0
9.12/0.57 = 16.0

In all cases the quotients are integers (to precisely so to be credible as experimental data), so it can be hypothesised that these quotients are the actual number of excess electrons on the drops, in which case the charge on an electron is 0.57*10^-14z.